How do I find the trigonometric function and the requested value?
1. Write the trigonometric function in terms of x
A) cos 0 if tan 0 = x and 0 are in quadrant II



2. Find the requested value
A) Find cot 0 if sec 0 = 25/7 and csc 0 = -25/24



B) Find cot 0 if tan 0 = √7/3 and 0 is in quadrant III

Answer by lwsshak3(11628)

How do I find the trigonometric function and the requested value?
1. Write the trigonometric function in terms of x
***
A) cos 0 if tan 0 = x and 0 are in quadrant II in which cos<0, sin>0,tan<0
tan 0=x/1=sin/cos
hypotenuse=√(x^2+1^2)=√(x^2+1)
cos 0=-1/√(x^2+1)
2. Find the requested value
A) Find cot 0 if sec 0 = 25/7 and csc 0 = -25/24
sec 0 = 25/7
cos 0=7/25
csc 0=-25/24
sin 0=-24/25
cot 0=cos/sin=-7/24

B) Find cot 0 if tan 0 = √7/3 and 0 is in quadrant III in which tan>0
cot 0=1/tan 0=3/√7=3√7/7